# Cluster-Dags as Powerful Background Knowledge For Causal Discovery

### v1

@@ 1 Introduction
Understanding causal relationships is essential for scientific inquiry and reasoning about the world. Researchers have always leveraged active experimentation and interventions to uncover causal mechanisms. But in many cases, such experiments are impractical or ethically off-limits—for instance, it would be unethical to expose communities to varying levels of environmental pollution to study their effect on health. The challenges of experimentation, along with the data explosion in recent decades, have led to continued methodological advances in causal discovery from observational data (guo2020survey; malinsky2018causal; peters2017elements).
One common approach to represent causal assumptions is the Structural Causal Model (SCM) framework (pearl2009causal; peters2017elements), which encodes these relationships in Directed Acyclic Graphs (DAGs), with edges representing the “directed flow of causal influence” between variables. One landmark result in causal discovery is that, without further assumptions, purely from observational data one can only identify the Markov equivalence class—a collection of graphs that are all consistent with the observed data. In large graphs, this ambiguity can make it difficult to pinpoint precise causal insights (spirtes2000causation; spirtes1995causal). Moreover, as the number of variables increases, the space of DAGs grows exponentially, posing serious scalability issues for many causal discovery algorithms (ijcai2024p374).
One way to help address these challenges is by incorporating background knowledge—partial information about the graph structure that can narrow the search space and guide causal inference. A recent method for encoding such partial knowledge is via Cluster-DAGs (C-DAG) (anand2023causal), which organize variables into clusters and assume that possible causal relationships between these clusters are known, but the causal structure within the clusters as well as the precise connections of individual variables across clusters is still unknown. By using structural information at this higher level, C-DAGs offer a way to manage complexity in high-dimensional settings. C-DAGs were originally developed for direct use in causal inference, reasoning about the strength of cause-effect relationships using the C-DAG directly (anand2023causal).
We leverage C-DAGs as background knowledge to improve constraint-based causal discovery methods using the prior knowledge during discovery process itself. We also extensively compare it to the related tiered background knowledge proposed by (andrews2020completeness). Our contributions include:
  • We show that tiered background knowledge (TBK) (andrews2020completeness) can be represented as a C-DAG, but not vice-versa—C-DAGs capture strictly more flexible types of background knowledge. In addition, C-DAGs can accommodate latent confounding between clusters, whereas TBK has to assume it non-existent, see Section˜2.1.
@@
  • C-DAGs allow warm-starting of constraint-based causal discovery algorithms, namely PC (spirtes2000causation) and FCI (zhang2008completeness) by pruning and orienting edges before running the discovery algorithm. This results in fewer CI tests needing to be performed. We introduce the Cluster-PC and Cluster-FCI formally and show their soundness (C-PC and C-FCI) and completeness (for C-PC only, C-FCI is not be complete by design).
@@ 1.1 Related work
There are two main ways in which background knowledge can be described in graphical structures: groupwise background knowledge and pairwise background knowledge. Pairwise background knowledge (fang2022representationcausalbackgroundknowledge; meek2013causalinferencecausalexplanation) restricts the relationship between pairs of variables, i.e., requiring or ruling out directed edges or ancestral relationships. In contrast, groupwise background knowledge (andrews2020completeness; brouillard2022typing; anand2023causal) organizes variables into different groups and then restricts the edges between these groups, without committing to any pairwise constraint directly.
@@
Previous work on background knowledge includes guiding score-based methods like KGS (hasan2024optimizingdatadrivencausaldiscovery), which can use prior knowledge on the absence/presence of an (un)directed edge, $A^{*}$-based methods (kleinegesse2022domainknowledgeabasedcausal), using absence/ presence of directed edges and tiers (non-ancestral constraints) and NOTEARS (chowdhury2023evaluationinducedexpertknowledge), using absence/ presence of directed edges). Most of these methods focus on pairwise background knowledge. For constraint-based methods, except for (andrews2020completeness), background knowledge is usually used to orient additional edges after receiving the CPDAG (brouillard2022typing; bang2023wisertimecausalequivalence) from the PC algorithm (spirtes2000causation). fang2022representationcausalbackgroundknowledge study the representation of causal background knowledge using pairwise background knowledge. Pairwise and groupwise background knowledge can be combined, although to the best of our knowledge, not much research has gone in this direction yet.
@@
Recently, different aspects of causal diagrams over groups of variables have been discussed. parviainen2017learning study learning groups of variables in Bayesian networks. wahl2023vector discuss two methods for inferring causal relationships between two groups of variables. melnychuk2024boundsrepresentationinducedconfoundingbias use C-DAGs to group confounders as a cluster. wahl2024foundations extend the theoretical framework of anand2023causal, discussing the relationships between micro and group level graphs. C-DAGs and other variable aggregation methods have increasingly been used to model causal systems (ma2025foundationmodelscausalinference; plecko2024mind; raghavan2025counterfactual; assaad2025identifiabilitytotaleffectsabstractions; xia2025causal; tabell2025clustering). Such an aggregate model can then, with our approach, be used to inform causal discovery of the more granular DAG.
@@
li2024disentangled present a method for learning disentangled causal representations from very high-dimensional data like images. Group graph variants have also increasingly become targets for causal discovery (niulearning; ninad2025causaldiscoveryvectorvaluedvariables; gobler2025nonlinearcausaldiscoverygrouped; anand2025causal). xia2024neural use neural causal models to learn neural causal abstractions by clustering variables and their domains, while massidda2024learningcausalabstractionslinear create a variant of LiNGAM (shimizu2014lingam), Abs-LiNGAM, to learn causal abstractions. li2025identifiabilitycausalabstractions; yvernes2025identifiabilitycausalabstractionshierarchy investigate the identifiability of causal abstractions, schooltink2025aligninggraphicalfunctionalcausal bridge graphical and functional causal abstractions and massidda2025weakly study causal sufficiency for causal abstractions. While much work focuses on learning the aggregated group graph itself, our approach uniquely leverages the C-DAG as assumed prior knowledge to perform causal discovery that resolves the relationships within the underlying micro-variables.
@@
The CaGreS algorithm by zeng2025causaldagsummarizationfull summarizes DAGs via node contractions and creates summary causal graphs (SCGs) with preserved utility for causal inference. ferreira2025identifyingC extend on anand2023causal by analyzing theoretical properties of SCGs, which are similar to C-DAGs, but also allow for cycles. yvernes2025completecharacterizationadjustmentsummary; assaad2025identifiabilitytotaleffectsabstractions; assaad2025identifiabilitymicrototaleffects; yvernes2025note; yvernes2025identifiabilitycommonbackdoorsummary; ferreira2025averagecontrolledaveragenatural investigate identifiability in SCGs, while ferreira2025identifyingmacrocausaleffects do so for cluster directed mixed graphs, a related concept. Transit clusters (tikka2023clustering), while similar, are specifically designed to cluster variables while preserving identifiability properties. zhu2024meaningful show that interventions in aggregation of variables (e.g., a surjective but non-injective function of cluster variables) are no longer well-defined. We build on these properties developed for C-DAGs to guide a constraint-based search for the detailed underlying DAG.
@@
On the application side, C-DAGs have also shown some promise for improving causal inference tasks. ribeiro2025bites apply causal inference to assess malaria risk, and they mention that causal discovery at the cluster level improves interpretability. anand2025leveraging apply C-DAGs to determine causal effects in medicine.
@@ 1.2 Preliminaries on causality and constraint-based causal discovery
We briefly introduce the most relevant preliminaries for structural causal models (SCM) and causal discovery. For more details, we refer the reader to (spirtes2000causation; pearl2009causality; peters2017elements). Throughout, we assume a fixed set of random variables $X_{1},\ldots,X_{n}$, which also serve as the node set of graphs $V=\{X_{1},\ldots,X_{n}\}$.
@@
For a one-to-one correspondence between d-separations in a graph $G=(V,E)$ and conditional independencies in a distribution $P(X_{1},\ldots,X_{n})$ to hold also requires the converse implication—called faithfulness. Faithfulness is not generally satisfied for the observational distribution and graph entailed by an SCM, but often assumed to hold in practice. The subtleties of the faithfulness assumption and its violations have been studied extensively in the causal discovery literature (zhang2002strong; ramsey2012adjacency; uhler2013geometry; marx2021weakerfaithfulnessassumptionbased).
Under the assumptions of causal sufficiency—there are no unobserved common causes of any variables in XX—and faithfulness, d-separations and conditional independencies are in one-to-one correspondence enabling constraint based causal discovery: by testing all possible conditional independencies in $P(X_{1},\ldots,X_{n})$ one can derive all d-separations that hold in $G$. It turns out that the set of DAGs that satisfies a given set of d-separations, called the Markov equivalence class, can be characterized as follows: Each DAG in the Markov equivalence class as the same skeleton (the same set of adjacencies) and the same set of v-structures (pearl2009causal; zhang2008causal). The Markov equivalence class can be represented by a partially directed graph (some directed, some undirected edges). If causal sufficiency is not assumed, the situation becomes more complicated and one can only infer more general types of graphs by testing conditional independencies. The two landmark algorithms for constraint based causal discovery with and without causal sufficiency are the PC and FCI algorithm, respectively. We will first focus on leveraging prior knowledge for a PC type algorithm in the fully observed case, before covering partial observations and an extension of FCI in Section˜3.2.
@@ 2 Cluster-DAGs
Cluster-DAGs (C-DAGs) were introduced by anand2023causal as a way to perform causal inference when one can not specify the entire DAG, but has enough information to organize groups of variables into a DAG. C-DAGs aggregate variables into clusters and then define macro-relationships between these clusters. An example can be seen in Fig.˜1.
@@
Many different DAGs may be compatible with the same C-DAG and vice versa, so C-DAGs form an equivalence class of DAGs. anand2023causal show that causal effects are still identifiable from the C-DAG $G_{C}$ if they are identifiable in every DAG $G$ compatible with $G_{C}$. Another important result is the soundness and completeness of d-separation in C-DAGs: any d-separation in the C-DAG also holds in any compatible DAG.
Theorem 1 (Soundness and completeness of d-separation in C-DAGs, (anand2023causal)).
@@
Based on these results, anand2023causal then develop an ID algorithm that is sound an complete for identifying causal effects in C-DAGs. In this work, we leverage implications of Theorem˜1 for causal discovery instead.
In constraint-based causal discovery (spirtes2000causation; zhang2008completeness), one typically assumes faithfulness and removes edges from an initially fully connected graph whenever a conditional independence is found between any pair of vertices. The key idea is that Theorem˜1 suggests that whenever $X\in C_{i},Y\in C_{j}$ and $C_{i}\perp\!\!\!\perp_{G_{C}}C_{j}\mid S$, then in any $G$ compatible with $G_{C}$ we have $X\perp\!\!\!\perp_{G}Y\mid S$ (slightly abusing notation in that $S$ is a set of clusters or a union over them, respectively) and thus $X,Y$ cannot be adjacent. Hence, C-DAGs allow us to aggressively prune a fully connected graph to warm-start constraint-based causal discovery on the micro-variables. Besides this warm start that potentially saves many conditional independence tests, the C-DAG structure can also be used to further speed up the remaining discovery process, which we discuss in Section˜3.
@@ 2.1 Comparison to tiered background knowledge
Definition 2 (Tiered background knowledge, (andrews2020completeness)).
@@
Similar to C-DAGs, TBK enforces the orientations of certain edges due to the macro-constraints. While causal discovery algorithms for TBK like FCI-Tiers (andrews2020completeness) have been developed, C-DAGs are strictly more flexible than TBK:
@@
    TBK cannot represent settings like the C-DAG $C_{1}\rightarrow C_{3}\leftarrow C_{2}$. TBK would have to put either $C_{1}$ or $C_{2}$ first in the tier ordering, which would only restrict the orientiation of edges between $C_{1},C_{2}$, but it cannot encode the strict absence of edges between $C_{1},C_{2}$ as the C-DAG does.
@@ 3.1 Cluster-PC
First, we consider C-DAGs without latent confounding, neither within nor between clusters. The assumed C-DAG structure over the micro variables then allows for immediate pruning of the intial complete graph and also for partial orientation of edges via Algorithm˜1. The resulting graph is an MPDAG (see Definition˜11), which can contain both directed and undirected edges. A detailed explanation can be found in Section˜A.1.
@@
The proof can be found in Section˜A.3. When a CI oracle is available, C-PC and PC with post-processing according to the C-DAG return the same result, so C-PC seemingly adds no value. However, in practice CI testing is an inherently difficult problem (ShahPeters2020hardness; Lundborg_2022) and a whole line of works investigates how the number of CI tests can be reduced to render constraint based causal discovery more effective (xie2008recursive; zhang2024membership; shiragur2024causal). In Section˜4 we demonstrate that avoiding unnecessary CI tests by incorporating the prior knowledge in C-PC indeed leads to substantial performance improvements.

### v2

@@ 1 Introduction
Understanding causal relationships is essential for scientific inquiry and reasoning about the world. Researchers have always leveraged active experimentation and interventions to uncover causal mechanisms. But in many cases, such experiments are impractical or ethically off-limits—for instance, it would be unethical to expose communities to varying levels of environmental pollution to study their effect on health. The challenges of experimentation, along with the data explosion in recent decades, have led to continued methodological advances in causal discovery from observational data (Guo et al., 2020; Malinsky & Danks, 2018; Peters et al., 2017).
One common approach to represent causal assumptions is the Structural Causal Model (SCM) framework (Pearl, 2009a; Peters et al., 2017), which encodes these relationships in Directed Acyclic Graphs (DAGs), with edges representing the “directed flow of causal influence” between variables. One landmark result in causal discovery is that, without further assumptions, purely from observational data one can only identify the Markov equivalence class—a collection of graphs that are all consistent with the observed data. In large graphs, this ambiguity can make it difficult to pinpoint precise causal insights (Spirtes et al., 2000; 1995). Moreover, as the number of variables increases, the space of DAGs grows exponentially, posing serious scalability issues for many causal discovery algorithms (Ganian et al., 2024).
One way to help address these challenges is by incorporating background knowledge—partial information about the graph structure that can narrow the search space and guide causal inference. A recent method for encoding such partial knowledge is via Cluster-DAGs (C-DAG) (Anand et al., 2023), which organize variables into clusters and assume that possible causal relationships between these clusters are known, but the causal structure within the clusters as well as the precise connections of individual variables across clusters is still unknown. By using structural information at this higher level, C-DAGs offer a way to manage complexity in high-dimensional settings. C-DAGs were originally developed for direct use in causal inference, reasoning about the strength of cause-effect relationships using the C-DAG directly (Anand et al., 2023).
We leverage C-DAGs as background knowledge to improve constraint-based causal discovery methods using the prior knowledge during discovery process itself. We also extensively compare it to the related tiered background knowledge proposed by (Andrews et al., 2020). Our contributions include:
  • We show that tiered background knowledge (TBK) (Andrews et al., 2020) can be represented as a C-DAG, but not vice-versa—C-DAGs capture strictly more flexible types of background knowledge. In addition, C-DAGs can accommodate latent confounding between clusters, whereas TBK has to assume it non-existent, see Section˜2.1.
@@
  • C-DAGs allow warm-starting of constraint-based causal discovery algorithms, namely PC (Spirtes et al., 2000) and FCI (Zhang, 2008b) by pruning and orienting edges before running the discovery algorithm. This results in fewer CI tests needing to be performed. We introduce the Cluster-PC and Cluster-FCI formally and show their soundness (C-PC and C-FCI) and completeness (for C-PC only, C-FCI is not be complete by design).
@@ 1.1 Related work
There are two main ways in which background knowledge can be described in graphical structures: groupwise background knowledge and pairwise background knowledge. Pairwise background knowledge (Fang et al., 2025; Meek, 1995) restricts the relationship between pairs of variables, i.e., requiring or ruling out directed edges or ancestral relationships. In contrast, groupwise background knowledge (Andrews et al., 2020; Brouillard et al., 2022; Anand et al., 2023) organizes variables into different groups and then restricts the edges between these groups, without committing to any pairwise constraint directly.
@@
Previous work on background knowledge includes guiding score-based methods like KGS (Hasan & Gani, 2024), which can use prior knowledge on the absence/presence of an (un)directed edge, $A^{*}$-based methods (Kleinegesse et al., 2022), using absence/ presence of directed edges and tiers (non-ancestral constraints) and NOTEARS (Chowdhury et al., 2023), using absence/ presence of directed edges). Most of these methods focus on pairwise background knowledge. For constraint-based methods, except for (Andrews et al., 2020), background knowledge is usually used to orient additional edges after receiving the CPDAG (Brouillard et al., 2022; Bang & Didelez, 2023) from the PC algorithm (Spirtes et al., 2000). Fang et al. (2025) study the representation of causal background knowledge using pairwise background knowledge. Pairwise and groupwise background knowledge can be combined, although to the best of our knowledge, not much research has gone in this direction yet.
@@
Recently, different aspects of causal diagrams over groups of variables have been discussed. Parviainen & Kaski (2017) study learning groups of variables in Bayesian networks. Wahl et al. (2023) discuss two methods for inferring causal relationships between two groups of variables. Melnychuk et al. (2024) use C-DAGs to group confounders as a cluster. Wahl et al. (2024) extend the theoretical framework of Anand et al. (2023), discussing the relationships between micro and group level graphs. C-DAGs and other variable aggregation methods have increasingly been used to model causal systems (Ma et al., 2025; Plecko & Bareinboim, 2024; Raghavan & Bareinboim, 2025; Assaad et al., 2024; Xia & Bareinboim, 2025; Tabell et al., 2025). Such an aggregate model can then, with our approach, be used to inform causal discovery of the more granular DAG.
@@
Li et al. (2024) present a method for learning disentangled causal representations from very high-dimensional data like images. Group graph variants have also increasingly become targets for causal discovery (Niu et al., 2022; Ninad et al., 2025; Göbler et al., 2025; Anand et al., 2025). Xia & Bareinboim (2024) use neural causal models to learn neural causal abstractions by clustering variables and their domains, while Massidda et al. (2024) create a variant of LiNGAM (Shimizu, 2014), Abs-LiNGAM, to learn causal abstractions. Li et al. (2025); Yvernes et al. (2025d) investigate the identifiability of causal abstractions, Schooltink & Zennaro (2025) bridge graphical and functional causal abstractions and Massidda et al. (2025) study causal sufficiency for causal abstractions. While much work focuses on learning the aggregated group graph itself, our approach uniquely leverages the C-DAG as assumed prior knowledge to perform causal discovery that resolves the relationships within the underlying micro-variables.
@@
The CaGreS algorithm by Zeng et al. (2025) summarizes DAGs via node contractions and creates summary causal graphs (SCGs) with preserved utility for causal inference. Ferreira & Assaad (2025b) extend on Anand et al. (2023) by analyzing theoretical properties of SCGs, which are similar to C-DAGs, but also allow for cycles. Yvernes et al. (2025a); Assaad et al. (2024); Assaad (2025); Yvernes et al. (2025b; c); Ferreira & Assaad (2025a) investigate identifiability in SCGs, while Ferreira & Assaad (2025c) do so for cluster directed mixed graphs, a related concept. Transit clusters (Tikka et al., 2023), while similar, are specifically designed to cluster variables while preserving identifiability properties. Zhu et al. (2024) show that interventions in aggregation of variables (e.g., a surjective but non-injective function of cluster variables) are no longer well-defined. We build on these properties developed for C-DAGs to guide a constraint-based search for the detailed underlying DAG.
@@
On the application side, C-DAGs have also shown some promise for improving causal inference tasks. Ribeiro et al. (2025) apply causal inference to assess malaria risk, and they mention that causal discovery at the cluster level improves interpretability. Anand & Hripcsak (2025) apply C-DAGs to determine causal effects in medicine.
@@ 1.2 Preliminaries on causality and constraint-based causal discovery
We briefly introduce the most relevant preliminaries for structural causal models (SCM) and causal discovery. For more details, we refer the reader to (Spirtes et al., 2000; Pearl, 2009b; Peters et al., 2017). Throughout, we assume a fixed set of random variables $X_{1},\ldots,X_{n}$, which also serve as the node set of graphs $V=\{X_{1},\ldots,X_{n}\}$.
@@
For a one-to-one correspondence between d-separations in a graph $G=(V,E)$ and conditional independencies in a distribution $P(X_{1},\ldots,X_{n})$ to hold also requires the converse implication—called faithfulness. Faithfulness is not generally satisfied for the observational distribution and graph entailed by an SCM, but often assumed to hold in practice. The subtleties of the faithfulness assumption and its violations have been studied extensively in the causal discovery literature (Zhang & Spirtes, 2002; Ramsey et al., 2006; Uhler et al., 2013; Marx et al., 2021).
Under the assumptions of causal sufficiency—there are no unobserved common causes of any variables in XX—and faithfulness, d-separations and conditional independencies are in one-to-one correspondence enabling constraint based causal discovery: by testing all possible conditional independencies in $P(X_{1},\ldots,X_{n})$ one can derive all d-separations that hold in $G$. It turns out that the set of DAGs that satisfies a given set of d-separations, called the Markov equivalence class, can be characterized as follows: Each DAG in the Markov equivalence class as the same skeleton (the same set of adjacencies) and the same set of v-structures (Pearl, 2009a; Zhang, 2008a). The Markov equivalence class can be represented by a partially directed graph (some directed, some undirected edges). If causal sufficiency is not assumed, the situation becomes more complicated and one can only infer more general types of graphs by testing conditional independencies. The two landmark algorithms for constraint based causal discovery with and without causal sufficiency are the PC and FCI algorithm, respectively. We will first focus on leveraging prior knowledge for a PC type algorithm in the fully observed case, before covering partial observations and an extension of FCI in Section˜3.2.
@@ 2 Cluster-DAGs
Cluster-DAGs (C-DAGs) were introduced by Anand et al. (2023) as a way to perform causal inference when one can not specify the entire DAG, but has enough information to organize groups of variables into a DAG. C-DAGs aggregate variables into clusters and then define macro-relationships between these clusters. An example can be seen in Fig.˜1.
@@
Many different DAGs may be compatible with the same C-DAG and vice versa, so C-DAGs form an equivalence class of DAGs. Anand et al. (2023) show that causal effects are still identifiable from the C-DAG $G_{C}$ if they are identifiable in every DAG $G$ compatible with $G_{C}$. Another important result is the soundness and completeness of d-separation in C-DAGs: any d-separation in the C-DAG also holds in any compatible DAG.
Theorem 1 (Soundness and completeness of d-separation in C-DAGs, (Anand et al., 2023)).
@@
Based on these results, Anand et al. (2023) then develop an ID algorithm that is sound an complete for identifying causal effects in C-DAGs. In this work, we leverage implications of Theorem˜1 for causal discovery instead.
In constraint-based causal discovery (Spirtes et al., 2000; Zhang, 2008b), one typically assumes faithfulness and removes edges from an initially fully connected graph whenever a conditional independence is found between any pair of vertices. The key idea is that Theorem˜1 suggests that whenever $X\in C_{i},Y\in C_{j}$ and $C_{i}\perp\!\!\!\perp_{G_{C}}C_{j}\mid S$, then in any $G$ compatible with $G_{C}$ we have $X\perp\!\!\!\perp_{G}Y\mid S$ (slightly abusing notation in that $S$ is a set of clusters or a union over them, respectively) and thus $X,Y$ cannot be adjacent. Hence, C-DAGs allow us to aggressively prune a fully connected graph to warm-start constraint-based causal discovery on the micro-variables. Besides this warm start that potentially saves many conditional independence tests, the C-DAG structure can also be used to further speed up the remaining discovery process, which we discuss in Section˜3.
@@ 2.1 Comparison to tiered background knowledge
Definition 2 (Tiered background knowledge, (Andrews et al., 2020)).
@@
Similar to C-DAGs, TBK enforces the orientations of certain edges due to the macro-constraints. While causal discovery algorithms for TBK like FCI-Tiers (Andrews et al., 2020) have been developed, C-DAGs are strictly more flexible than TBK:
@@
    TBK cannot represent settings like the C-DAG $C_{1}\rightarrow C_{3}\leftarrow C_{2}$. TBK would have to put either $C_{1}$ or $C_{2}$ first in the tier ordering, which would only restrict the orientation of edges between $C_{1},C_{2}$, but it cannot encode the strict absence of edges between $C_{1},C_{2}$ as the C-DAG does.
@@ 3.1 Cluster-PC
First, we consider C-DAGs without latent confounding, neither within nor between clusters. The assumed C-DAG structure over the micro variables then allows for immediate pruning of the initial complete graph and also for partial orientation of edges via Algorithm˜1. The resulting graph is an MPDAG (see Definition˜11), which can contain both directed and undirected edges. A detailed explanation can be found in Section˜A.1.
@@
The proof can be found in Section˜A.3. When a CI oracle is available, C-PC and PC with post-processing according to the C-DAG return the same result, so C-PC seemingly adds no value. However, in practice CI testing is an inherently difficult problem (Shah & Peters, 2020; Lundborg et al., 2022) and a whole line of works investigates how the number of CI tests can be reduced to render constraint based causal discovery more effective (Xie & Geng, 2008; Zhang et al., 2024; Shiragur et al., 2024). In Section˜4 we demonstrate that avoiding unnecessary CI tests by incorporating the prior knowledge in C-PC indeed leads to substantial performance improvements.